Utilities $1,000

Rent    $8,000

asistant Salary $5,000

managers salary $6000

Fixed cost =20,000

coco butter per box of chocolate $2

milk per $o.40

sugar per $1.60

secret ingredient $3

+Wages of the employee $3

= variable cost is 10

selling price is = $20

  1. Calculate break-even quantity
  2. What is the value in dollars of the number of Hershey's that they need to break even?
  3. the market share that they need to capture breakeven?
Answer #1

Given, Fixed cost = $20,000

Pricing per unit = $20

Variable cost per unit = $10

a)

Break even quantity = Fixed cost/(pricing per unit – variable cost per unit)

Break even quantity = 20,000/(20-10)

Break even quantity =2000

b)

value in dollars = Fixed cost + total variable cost = total pricing

value in dollars = 20000+(10*2000)=20*2000= 40000

c)

suppose, there is monthly demand of 20000 and breakeven point or quantity 2000, hence market share would be 2000/20000 = 10%, that can be achievable